OpenStudy (anonymous):

Solve x - y/3 = 2 for y.

OpenStudy (alexandervonhumboldt2):

first subtract x from both sides: x-x-y/3=2-x -y/3=2-x multiply both sides by -3. y=(2-x)*(-3)=-6+3x

OpenStudy (anonymous):

thank you:)))

OpenStudy (alexandervonhumboldt2):